OpenStudy (anonymous):

A ball is dropped from the top of a 349-foot building. How many seconds will it take to hit the ground if the height is given by h = −16t2+ s, where s is the initial height? Answer in units of sec

OpenStudy (jlastino):

Just input 349 as h then solve for s

OpenStudy (jlastino):

oh wait aren't the variable's switched??

OpenStudy (anonymous):

I don't know i dont understand this?? problem

OpenStudy (lgbasallote):

you're looking for t so isolate it... h - s = -16t^2 t^2 = \(\large \frac{(h-s)}{-16}\) t = \(\large \sqrt{\frac{(h-s)}{16}}\) input your values... hint: h = 349 ft s = 0 ft

OpenStudy (anonymous):

what is the height of the ball on the ground? [BIG hint]

OpenStudy (lgbasallote):

wait....the denom in the last line should be negative...asd in the hint s = 349 h = 0 sorry

OpenStudy (anonymous):

thanks i got it know :) God bless